OpenStudy (maddytarin_98):

Find the inverse of each of the given functions. f(x) = 4x − 12

sam (.sam.):

To find the inverse function, let y=f(x), then solve for 'x', the 'x' later will be your inverse function, \[f(x) = 4x − 12\] \[y=4x-12\] \[y+12=4x\] \[\frac{y+12}{4}=x\] \[f^{-1}(x)=\frac{x+12}{4}\]